OpenStudy (anonymous):

Identify the y-intercept of the line below. 5x - 8y = -120

OpenStudy (anonymous):

15

Directrix (directrix):

When x = 0, the y-intercept can be found. 5x - 8y = -120 -8y = -120 y = 15